# Data Integration with Fusion Searchlight: Classifying Brain States from Resting-state fMRI

### v1

@@ 4 Discussion
In our study we presented an interpretable SL framework for fusing the complementary information of different rs-fMRI metrics. We first used an artificial dataset to study how different combinations of informative and uninformative data sources can affect decoding capabilities of FuSL. We demonstrated that across different signal to noise ratios the gain of adding informative sources exceeded the loss in performance when adding an uninformative source (2). We further illustrated how reconstructing feature-importance using SHAP can recover the impact of each data source, as well as the exact location of informative vertices. In a next step we studied how combining ReHo, fALFF and FCE can enhance the prediction of pharmacological treatment in rs-fMRI. Decoding based on combinations of different rs-fMRI metrics outperformed consistently decoding based on single metrics (figure 4). A combination of measuring localized fluctuations in low-frequency fluctuations and global connectivity efficiency (fALFF||||| |FCE), as well as the combination of all three metrics (ReHo||||| |fALFF||||| |FCE), achieved highest decoding performances. Especially in clinical studies where collecting a very large sample size is often not feasible, the integration of different information sources can help improving decoding sensitivity on smaller datasets. Still, how distinct rs-fMRI markers are coupled and related to each other is not yet investigated and understood in detail (Lv et al., 2018). Our comparisons show that metrics derived from the same signal can exhibit very distinct characteristics in different areas across the cortex. For instance in the somatomotor network we observed an increase both in fALFF and ReHo, but within the visual network we observed an increase in fALFF whereas ReHo decreased. Note that the interrelations of rs-fMRI markers and their optimal combination in FuSL might therefore also vary between different rs-fMRI studies. It has yet to be shown which of the patterns reported here will remain stable across different studies. Furthermore, we have shown how one can utilize explainable AI to reconstruct the impact of each rs-fMRI metric on the decoding in our workflow. Incorporating the SHAP framework (Lundberg and Lee, 2017) allowed us to analyze locally the importance of features (in our study rs-fMRI metrics), which we combined with local informativeness (derived from decoding accuracy) to quantify the impact of a feature across the cortex. We propose to additionally incorporate local differences in features between groups to visualize whether a feature-expression increases or decreases. In theory a large feature difference does not necessarily correlate with a high feature impact, because the output of a non-linear decoding model can depend on a highly non-linear combination of its input features. In our practical application we however observed that absolute amplitudes of the impact and feature-weighted impact corresponded qualitatively very well (supplementary figure S3), so for a compact interpretation we considered it useful to visualize the feature-weighted impact. Yet, to study the impact of a feature in more detail, we propose to additionally visualize the impact in isolation. Besides using SHAP for computing the differential impact of each metric, we have shown that incorporating explainable AI can be generally beneficial for searchlight analysis. Evaluating feature-importance of the decoding model could help to reconstruct the exact location of informative vertices or voxels, which makes this technique interesting for all types of searchlight decoding analysis.
@@
In our study we showed how fusing different metrics can improve SL decoding specifically in resting-state fMRI analysis. Our framework thereby leverages the strengths of SL decoding, inherently limiting model complexity by using spatially localized features, making it a robust tool for neuroimaging studies with smaller sample sizes. Furthermore using a model-agnostic approach for measuring feature importance based on Shapley values (Lundberg and Lee, 2017), our workflow allows one to integrate any kind of decoding model. Besides this application of decoding in rs-fMRI, our FuSL workflow can be easily adapted to various additional domains of neuroimaging. Combining different imaging modalities in FuSL could be beneficial for studying processes and conditions that simultaneously affect brain structure and function like observed in learning (Carreiras et al., 2009) or schizophrenia (Dabiri et al., 2022). Also this framework would allow one to integrate activation maps of combined fMRI and EGG studies (Menon and Crottaz‐Herbette, 2005). In this manner FuSL provides a flexible and interpretable tool for diverse applications in neuroimaging analysis.
@@
This work has been supported by the German Research Foundation (Deutsche Forschungsgemeinschaft) (DFG), project number 403161218, to JS and within the framework of FOR2858.
@@
with the explanation model $g(z^{\prime})$ and coalition vector $z^{\prime}\in\{0,1\}^{M}$, indicating whether a feature is present or not, $M$ the number of input features and Shapley values $\phi_{i}\in\mathbb{R}$ of feature $i$. These values can be approximated using different model specific and model agnostic methods described by Lundberg and Lee (Lundberg and Lee, 2017), whereby in our implementation we rely on the model-agnostic approach employing successive permutations of input features. We then compute the importance of an input feature $i$ as the average of absolute Shapley values of all $J$ test samples:
@@
Figure S1: Decoding performance of FuSL in dependence of different noise levels. (A) Decoding accuracies when using sources with higher signal to noise ratios ($S/N=2$). (A1) and (A2) show decoding accuracies for all combinations of sources in ROI 1 and 2 respectively, and (A3) the overall decoding accuracies averaged across ROI 1 and 2. (B) Decoding accuracies when using sources with lower signal to noise ratios ($S/N=0.1$). Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
@@
ReHo||||| |fALFF||||| |FCE vs. ReHo 0.0005 0.765
ReHo||||| |fALFF||||| |FCE vs. FCE 0.0068 0.600
ReHo||||| |fALFF||||| |FCE vs. fALFF 0.0881 0.509
ReHo||||| |fALFF||||| |FCE vs. ReHo||||| |FCE 0.0078 0.505
ReHo||||| |fALFF||||| |FCE vs. ReHo||||| |fALFF 0.0318 0.225
ReHo||||| |fALFF||||| |FCE vs. ReHo||||| |fALFF 0.2267 0.028
Figure S2: A comparison of significant vertices of combination ReHo||||| |fALFF||||| |FCE and all other combinations overlayed with resting-state networks defined by Yeo et al. (Yeo et al., 2011). (A) Locations where only ReHo||||| |fALFF||||| |FCE yields significant decoding accuracies are highlighted in red in the left (L) and right (R) cortex. Locations where both combinations can significantly decode are marked in white, and locations where only the other combination is able to decode are depicted in blue. (B) Displays the 7 resting-state networks defined by Yeo et al. (Yeo et al., 2011).
@@
Figure S4: Decoding accuracy and feature-weighted impact of FuSL overlayed with resting-state networks defined by Yeo et al. (Yeo et al., 2011). (A) Decoding accuracy of FuSL across the cortex based on a combination of ReHo, fALFF and FCE. (B) The feature-weighted impact allows us to identify which brain areas are informative and important for the decoding and whether ReHo, fALFF and FCE are increased or decreased due to the treatment of alprazolam. (C) Displays the 7 resting-state networks defined by Yeo et al. (Yeo et al., 2011).

### v2

@@ 4 Discussion
In our study we presented an interpretable SL framework for fusing the complementary information of different rs-fMRI metrics. We first used an artificial dataset to investigate how different combinations of informative and uninformative data sources can affect decoding capabilities of FuSL. We demonstrated that across a variety of different artificial datasets the gain of adding informative sources exceeded the loss in performance when adding an uninformative source. We further illustrated how reconstructing feature-importance using SHAP allowed us to recover the impact of each data source, as well as the exact location of informative vertices. In a next step we studied how combining ReHo, fALFF and FCE can enhance the prediction of pharmacological treatment in rs-fMRI. In this application combining fALFF with connectivity-based metrics (ReHo, FCE or both) could considerably enhance the decoding performance. Especially in clinical studies where collecting a very large sample size is often not feasible, the integration of different information sources can help improving decoding sensitivity on smaller datasets. Still, how distinct rs-fMRI markers are coupled and related to each other is not yet investigated and understood in detail (Lv et al., 2018). Our comparisons show that metrics derived from the same signal can exhibit very distinct characteristics in different areas across the cortex. For instance in the somatomotor network we observed an increase both in fALFF and ReHo, but within the visual network we observed an increase in fALFF whereas ReHo decreased. Note that the interrelations of rs-fMRI markers and their optimal combination in FuSL might therefore also vary between different rs-fMRI studies. It has yet to be shown which of the patterns reported here will remain stable across different studies.
In a next step we have shown how one can utilize explainable AI to reconstruct the impact of each rs-fMRI metric on the decoding in our workflow. Incorporating the SHAP framework (Lundberg and Lee, 2017) allowed us to analyze locally the importance of features (in our study rs-fMRI metrics), which we combined with local informativeness (derived from decoding accuracy) to quantify the impact of a feature across the cortex. We propose to additionally incorporate local differences in features between groups to visualize whether a feature-expression increases or decreases. In theory a large feature difference does not necessarily correlate with a high feature impact, because the output of a non-linear decoding model can depend on a highly non-linear combination of its input features. In our practical application we however observed that absolute amplitudes of the impact and feature-weighted impact corresponded qualitatively very well (supplementary figure S10), so for a compact interpretation we considered it useful to visualize the feature-weighted impact. Yet, to study the impact of a feature in more detail, we propose to additionally visualize the impact in isolation. Besides using SHAP for computing the differential impact of each metric, we have shown that incorporating explainable AI can be generally beneficial for searchlight analysis. Evaluating feature-importance of the decoding model could help to reconstruct the exact location of informative vertices or voxels, which makes this technique interesting for all types of searchlight decoding analysis.
@@
However, at this point the interpretation of above inconsistent directional changes of ReHo in different brain areas need to be regarded with some caution and seeks deeper investigations. Furthermore, the detailed mechanistic connections between pharmacological interventions and modulations in BOLD dynamics are relatively poorly understood to date, so also inferences from alterations in rs-fMRI markers to behavioral changes should consider these current knowledge gaps (Khalili-Mahani et al., 2017). Within the scope of our study we investigated the application of FuSL to rs-fMRI using only one dataset, so finding reliable and optimal combinations of rs-fMRI metrics likely requires more extensive research including potentially larger datasets. Furthermore, to what extent combining information from very distinct neuroimaging modalities improves decoding capabilities of FuSL has to be shown in future research.
Until now our study could provide first evidence for the potential of FuSL in MVPA decoding applications. While deep learning frameworks often require a large amount of training data and manual fine-tuning of model parameters (Schulz et al., 2020), SL has proven to be a robust and widely-used MVPA tool that limits its complexity inherently by using spatially localized information. Furthermore, by using a model-agnostic approach for measuring feature importance based on Shapley values (Lundberg and Lee, 2017), our workflow allows one to integrate any kind of MVPA decoding model. Besides this application of decoding in rs-fMRI, our FuSL workflow can be easily adapted to various additional domains of neuroimaging. Combining different imaging modalities in FuSL could be beneficial for studying processes and conditions that simultaneously affect brain structure and function like observed in learning (Carreiras et al., 2009) or schizophrenia (Dabiri et al., 2022). Also this framework would allow one to integrate activation maps of combined fMRI and EGG studies (Menon and Crottaz‐Herbette, 2005). In this manner FuSL provides a flexible and interpretable tool for diverse applications in neuroimaging analysis.
@@
This work has been supported by the German Research Foundation (Deutsche Forschungsgemeinschaft) (DFG), project number 403161218, to JS within the framework of FOR2858.
@@
The centrality of a node $i$ in a graph $G$ can be described by betweenness centrality, defined as the sum of the fraction of shortest paths between two nodes $j$ and $k$ that pass through node $i$:
$c_{B}(i)=\sum_{j,k\in G}\frac{\sigma(j,k|i)}{\sigma(j,k)}$ (5)
where $\sigma(j,k)$ is the number of shortest paths between $j$ and $k$, and $\sigma(j,k|i)$ the number of shortest paths passing through node $i$ (other than $j,k$).
@@
with the explanation model $g(z^{\prime})$ and coalition vector $z^{\prime}\in\{0,1\}^{M}$, indicating whether a feature is present or not, $M$ the number of input features and Shapley values $\phi_{i}\in\mathbb{R}$ of feature $i$. These values can be approximated using different model specific and model agnostic methods described by Lundberg and Lee (2017), whereby in our implementation we rely on the model-agnostic approach employing successive permutations of input features. We additionally compare this permutation based method for approximating Shapley values to linear explanation models (KernelSHAP). We then compute the importance of an input feature $i$ as the average of absolute Shapley values of all $J$ test samples:
@@
Figure S1: Spatial variability in the mesh density of the fsLR standard space mesh. (A) Distribution of number of vertices in a spherical SLs defined in the worldspace across the cortex. A larger radius (in $mm$) increases the standard deviation of number of vertices $\sigma$ within spherical SLs. (B) Average distance of vertices to SL center when using a constant number of k-hop neighbors. The variability in the size of the SLs across the cortex increases with a larger number of k-hop neighbors included in SLs.
Figure S2: Impact of MVPA model and search light (SL) radius on the decoding using artificial data. (A) A support vector classifier (SVC) with radial basis function kernel (RBF) outperforms SVC with a linear (lin) and polynomial (poly) kernel. Further it shows better decoding accuracy than a k-nearest neighbor (KNN) classifier, a random forest (RF) classifier based on 50 decision trees, and a ridge regression classifier. (B) A SL radius of order 3 provides a reasonable trade-off between accuracy and computation time. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
Figure S3: Decoding performance of FuSL on an artificial dataset with a signal standard deviation of $\sigma=1$ across samples, in dependence of different signal to noise ratios and different levels of spatial autocorrelations simulated by Gaussian smoothing. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
Figure S4: Decoding performance of FuSL on an artificial dataset with a signal standard deviation of $\sigma=2$ across samples, in dependence of different signal to noise ratios and different levels of spatial autocorrelations simulated by Gaussian smoothing. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
@@
ReHo||||| |fALFF||||| |FCE vs. FCE 0.0053 0.817
ReHo||||| |fALFF||||| |FCE vs. ReHo 0.0004 1.015
ReHo||||| |fALFF||||| |FCE vs. fALFF 0.0256 0.659
ReHo||||| |fALFF||||| |FCE vs. ReHo||||| |FCE 0.0127 0.634
ReHo||||| |fALFF||||| |FCE vs. fALFF||||| |FCE 0.0091 0.145
ReHo||||| |fALFF||||| |FCE vs. ReHo||||| |fALFF 0.1998 0.107
Figure S5: A comparison of decoding accuracy and significant vertices of combination ReHo||||| |fALFF||||| |FCE and all other combinations overlayed with resting-state networks defined by Yeo et al. (2011). (A) Local differences in decoding test accuracies between ReHo||||| |fALFF||||| |FCE and all other combinations in the left (L) and right (R) cortex. (B) Locations where only ReHo||||| |fALFF||||| |FCE yields significant decoding accuracies are highlighted in red. Locations where both combinations can significantly decode are marked in white, and locations where only the other combination is able to decode are depicted in blue. (C) Displays the 7 resting-state networks defined by Yeo et al. (2011).
Figure S6: Impact of MVPA model, search light (SL) radius and resting-state fMRI metric on the decoding of pharmacological treatment. (A) A support vector classifier (SVC) with radial basis function kernel (RBF) outperforms SVC with a linear (lin) and polynomial (poly) kernel. Further it shows better decoding accuracy than a k-nearest neighbor (KNN) classifier, a random forest (RF) classifier based on 50 decision trees, and a ridge regression classifier. (B) A SL radius of order 3 shows a reasonable trade-off between accuracy and computation time. (C) When combining different rs-fMRI metrics, fALLF shows to be slightly more informative than ALFF, FC efficiency (FCE) slightly more informative than FC betweeness centrality (FCbC), and ReHo with a k-hop neighborhood order of 5 is slightly more informative than of order 3. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ****: $p\leq 0.0001$, ns: not significant, (ns): not significant after false discovery rate correction.
Figure S7: Impact of within subject demeaning on the decoding accuracy. Removing the mean of rs-fMRI metrics within each subject removes the baseline variability of rs-fMRI metrics between different subjects and helps to better detect the effect of the alprazolam treatment, resulting inconsiderably higher decoding accuracies. Error bars represent 95% confidence intervals across folds. Significant differences of accuracies are indicated with: *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$, ns: not significant, (ns): not significant after false discovery rate correction.
Figure S8: Sensitivity, specificity and area under receiver operating characteristic (ROC) curve based on the decoding of the alprazolam treatment. Patterns of sensitivity (A), specificity (B) and area under ROC curve (C) resemble closely the pattern of decoding accuracy.
Figure S9: Comparison of methods to compute SHAP values for determining the impact of each rs-fMRI metric. (A) Impact based on SHAP values derived from permutation testing. (B) Impact based on SHAP values derived from Kernel SHAP. Both explanation approaches produce similar impact maps, whereby Kernel SHAP attributes slightly less importance to fALFF, than the permutation based approach. The computation time of the permutation based method is with 5.29 h lower than the time of the Kernel SHAP method requiring 17.97 h on a Intel(R) Xeon(R) Platinum 8268 CPU, running under Ubuntu 20.04 with 256 GB of RAM.
@@
Figure S11: Decoding accuracy and feature-weighted impact of FuSL overlayed with resting-state networks defined by Yeo et al. (2011). (A) Decoding accuracy of FuSL across the cortex based on a combination of ReHo, fALFF and FCE. (B) The feature-weighted impact allows us to identify which brain areas are informative and important for the decoding and whether ReHo, fALFF and FCE are increased or decreased due to the treatment of alprazolam. (C) Displays the 7 resting-state networks defined by Yeo et al. (2011).